A monopolist faces demand Q = 60 – P, and has a TC of Q + 2Q2. Then to maximize profit, the monopolist should produce ____ units of Q.

Group of answer choices

A. 9.83

B. 12.25

C. 14.11

D. 16.72
Answer #1

for propt maximization in Monopoly MC=AR) Here, TCa 2020 Mc = d(JC) do = d(20 2to) dQ The =40+1 Pa Gora 2. TR = (60-Q) xQ -